OpenStudy (anonymous):

The addition Property of Equality 1. for each of these equation *solve the equation *justify each step with either the addition property of equality, a number property or an operation *check the solution (write how you get the answer) a. x - 7 = 0 b.2x - 7 =x c. 6x - 7 =5x can someone solve this for me???

OpenStudy (hba):

Lemme show you some examples 2x+5=3 Just take the 3 on the other hand it is +3 when it crosses = sign it will become -3 2x+5-3=0 2x+2=0 If we shift +2 on the other hand it will become -2 2x=-2 and x=-1 Similarly x-5=0 Take the -5 on the other hand it will become +5 x=5 I hope you can do this yourself now.

OpenStudy (anonymous):

i.e. note that addition property of equality means that if x-a=0, where a is a number, then if you add a to it, you must add to both sides. x-a +a =0+a